40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 2154     Entry time: Wed Oct 28 05:02:28 2009
Author: rob 
Type: Update 
Category: Locking 
Subject: back 

LockAcq is back on track, with the full script working well.  Measurements in progress.

ELOG V3.1.3-